OpenStudy (anonymous):

Last question then I wont bother you nice people. Find the range of the function g(x) = 3x - xsquared with domain D={-1,0,1,2,3}

OpenStudy (chriss):

The range will be the set of answers you get from plugging the domain values into g(x). For instance the first value you get is from plugging -1 into the function. \[g(x) = 3x-x^{2}\] \[g(-1) = 3(-1)-(-1)^{2} = (-3)-1 = -4\] So the first element of your range will be -4.